OpenStudy (anonymous):

find the curve of intersection of the elliptic cyclinder x^2 +4z^2 = 16 and the plane x+3y-2z = 3

OpenStudy (anonymous):

One thing you could possibly do is multiply the plane by (16/3) giving: (16/3)x+16y-(32/3)z=16 Now they are both=16 x^2+4z^2=(16/3)x+16y-(32/3)z Then solve it however you want it. You could also solve the first equation for z or x but the problem is you have a sqrd so you end up with a +/- ambiguity.

OpenStudy (anonymous):

Parametrize the curve of intersection. The elliptic cylinder can be parametrized by x=4cos(t) z=2sin(t) (y=y) Then rewrite the equation of the plane as y=1+2/3z-1/3x, so y=1+4/3sin(t)-4/3cos(t) The curve is then parametrized by x=4cos(t) y=1+4/3sin(t)-4/3cos(t) z=2sin(t) t ranges from 0 to 2pi